# Contrastive Language–Structure Pre-training Driven by Materials Science Literature

### v1

@@ Acknowledgments
R. I. is partly supported by JSPS KAKENHI Grant Number 24K23910. N. C. is partly supported by JSPS KAKENHI Grant Number 21K14130. Y. U. is partly supported by JST-Mirai Program Grant Number JPMJMI21G2 and JST Moonshot R&D Program Grant Number JPMJMS2236.
Data availability
This study used the Crystallography Open Database (COD) [22] as the source of crystal structure data. Please see Sec. 7.27.4 for the data retrieval and processing details.
Code availability
We will release our code upon acceptance of this work for publication.
@@
YS conceived the concept, developed the model, conducted the analysis, and drafted the manuscript. TT provided guidance on the methodology, analysis and writing, assisted with the literature review, and revised the manuscript draft. RI and KS contributed expertise in materials science, advising on the analysis. NC contributed expertise in machine learning, advising on the methodology. YU co-led materials-related collaborations, provided guidance on the plan and methodology from a machine learning perspective, and advised on the writing. KO co-led materials-related collaborations and provided guidance on the plan and analysis from a materials science perspective.

### v2

@@ Acknowledgments
T.T. is partly supported by JSPS KAKENHI Grant Number 24K23911. R.I. is partly supported by JSPS KAKENHI Grant Number 24K23910. N.C. is partly supported by JSPS KAKENHI Grant Number 21K14130. Y.U. is partly supported by JST-Mirai Program Grant Number JPMJMI21G2 and JST Moonshot R&D Program Grant Number JPMJMS2236. K.O. is partly supported by the MEXT Program: Data Creation and Utilization-Type Material Research and Development Project (Digital Transformation Initiative Center for Magnetic Materials) Grant Number JPMXP1122715503.
Data availability statement
This study used the Crystallography Open Database (COD) [23] as the source of crystal structure data. Please see Sec. 7.27.4 for the data retrieval and processing details. All code for dataset creation, model implementation, training, and analysis, along with pretrained model weights and the resulting datasets, is available at: https://github.com/Toyota/clasp .
@@
Y.S. conceived the concept, developed the model, conducted the analysis, and drafted both the initial manuscript and the revision. T.T. provided guidance on the methodology, analysis and writing; assisted with the literature review; revised the manuscript draft; and led the review response and revision. R.I. and K.S. contributed expertise in materials science, advising on the analysis. N.C. contributed expertise in machine learning, advising on the methodology. Y.U. co-led materials-related collaborations, provided guidance on the plan and methodology from a machine learning perspective, and advised on the writing. K.O. co-led materials-related collaborations and provided guidance on the plan and analysis from a materials science perspective.